# Informed Asymmetric Actor-Critic: Leveraging Privileged Signals Beyond Full-State Access

### v1

@@ 6 EXPERIMENTS
We evaluate the proposed informed asymmetric actor-critic framework on POMDP benchmarks against three baseline actor-critic variants. Moreover, we empirically validate the effectiveness of the presented informativeness criteria on synthetic informed POMDP instances.
@@
6.1 Environments
We use six benchmark navigation tasks from the work of Baisero and Amato [10] to evaluate the learning performance of our method: Heaven-Hell-3, Shopping-5, Car-Flag, Cleaner, Memory-Four-Rooms-7x7, and Memory-Four-Rooms-9x9. Each environment is formulated as a POMDP, and we define task-specific privileged partial information accessible only to the critic. For the first two tasks, the privileged input corresponds to the Earth Mover’s Distance between the agent’s position and the target. In Car-Flag, the agent’s velocity is provided as additional information. For the remaining three environments, the privileged signal consists of an expanded spatial observation of the agent’s surroundings. See Appendix B for details.
To assess the expressiveness of our informativeness criteria, we generate synthetic informed POMDP instances with a finite state space ($|\mathcal{S}|=10$), a discrete action space ($|\mathcal{A}|=4$), and continuous observation and information spaces. Following the methodology of François-Lavet et al. [27], transition probabilities are randomly assigned by setting each $(s,a,s^{\prime})$-entry to zero with probability $0.75$, and sampling uniformly from $[0,1]$ otherwise. To ensure valid transitions, we assign a non-zero probability to a randomly chosen next state whenever all transitions from a given state-action pair are initially zero. We then normalize the probabilities to ensure they sum to one. Rewards are sampled uniformly from [1,1][-1,1] at initialization. Privileged information is generated by sampling from a Gaussian distribution centered on a state-specific embedding, with variance controlled by a noise parameter $\varsigma\in\mathbb{R}_{\geq 0}$. Observations are then obtained by applying a noisy linear transformation to the privileged information, with another noise parameter controlling the observation uncertainty.
6.2 Baselines
We compare our informed asymmetric actor-critic method against three advantage actor-critic (A2C) variants: (1) A2C, a symmetric approach using a history-based critic $\hat{V}(h)$; (2) asym-A2C-s, an asymmetric variant with a state-based critic $\hat{V}(s)$; and (3) asym-A2C-hs, an asymmetric variant with a history-state critic $\hat{V}(h,s)$. For the benchmark environments and all baselines, we adopt the model architectures and hyperparameters recommended by Baisero and Amato [10].
6.3 Results and Discussion
We highlight three key results from our evaluation: (a) empirical learning curve comparison; (b) results from the hypothesis test assessing our HSCIC-based informativeness criterion; and (c) the boxplot distribution of the test statistic for our post-hoc informativeness criterion.
Learning curves.
@@ 6.3 Results and Discussion
First, we compare informed-asym-A2C against the baselines across the six benchmark navigation tasks (Figure 1 (a)–(f)). In Heaven-Hell-3, informed-asym-A2C exhibits strong performance relative to A2C and asym-A2C-s, though it is slightly outperformed by asym-A2C-hs, which benefits from full-state access in the critic. In Shopping-5, the informed asymmetric actor-critic converges faster than asym-A2C-hs and achieves comparable final returns. In Car-Flag, informed-asym-A2C outperforms all baselines, demonstrating both higher sample efficiency and improved final performance. For Cleaner, asym-A2C-hs achieves marginally higher returns, but informed-asym-A2C converges at a similar rate with greater stability than A2C, which suffers a performance drop after 2.5 million steps. In the Memory-Four-Rooms tasks, informed-asym-A2C significantly outperforms both asymmetric baselines, particularly asym-A2C-s, which lacks critic access to history.
Overall, informed-asym-A2C matches or surpasses the performance of asym-A2C-hs and/or achieves faster convergence in most tasks while relying on less privileged information. This underscores the importance of not only leveraging asymmetric information but also structuring it to align effectively with task requirements.
Information $\bar{\mathscr{I}}$ (mean $\pm$ std) $p$-val. (mean $\pm$ std)
$i_{t}=\emptyset$ $66.287\pm 29.840$ $1.000\pm 0.000$
$i_{t}$: $\varsigma=0.0$ $132.435\pm 90.216$ $0.108\pm 0.144$
$i_{t}$: $\varsigma=0.1$ $140.454\pm 78.242$ $0.147\pm 0.194$
$i_{t}$: $\varsigma=0.5$ $122.671\pm 56.914$ $0.136\pm 0.229$
$i_{t}$: $\varsigma=0.9$ $98.173\pm 43.427$ $0.239\pm 0.235$
$i_{t}=s_{t}$ $126.793\pm 63.656$ $0.122\pm 0.251$
Table 1: HSCIC- and $p$-value statistics for different privileged signals, computed across 12 synthetic informed POMDP instances.
Kernel-based independence criterion.
To investigate the role and informativeness of different privileged signals, we estimate the empirical mean pointwise HSCIC value $\bar{\mathscr{I}}$ and its null distribution across 12 synthetic informed POMDP instances. Gaussian RBF kernels $k_{G}$, $k_{I}$, and $k_{Z}$ are used with bandwidths selected via the median heuristic, i.e., using the median pairwise distance between samples. We consider three configurations: (1) no privileged input, (2) noisy latent vectors $i_{t}$ with varying noise levels $\varsigma$, and (3) full-state access ($i_{t}=s_{t}$). Using random policies, we collect 20 episodes of length 25 and perform $B=30$ permutations.
Table 1 reports the mean and standard deviation of $\bar{\mathscr{I}}$ and the corresponding $p$-values across informed POMDP instances. As noise increases, both the mean and variance of $p$-values generally increase (except for $\varsigma=0.5$), suggesting reduced statistical significance and potential informativeness. Interestingly, full-state access appears on average less informative than noiseless partial input, consistent with trends observed in the learning curve analysis. However, the substantial variance across instances highlights the instance-specific nature of informativeness.
Figure 2: Boxplot distributions of $\epsilon$ over 1,000 test episodes for synthetic POMDP instances with privileged signal $i_{t}$ ($\varsigma=0.1$), computed for (a) different $\delta$ across 20 instances; (b) fixed $\delta=0.05$ for five randomly sampled instances.
$\mathbf{(\epsilon,\delta)}$-informativeness.
We validate the post-hoc informativeness criterion on $20$ synthetic informed POMDP instances with noisy privileged signals ($\varsigma=0.1$). Specifically, we compare critic variants trained for 25,000 episodes with $T=50$ using the Rec-NAC algorithm [14], with an Elman-type RNN of width 64 to encode the observation-action history. The RNN is followed by a fully connected layer with 256 units with ReLU activation, and a linear readout.
Figure 2 shows the boxplot distributions of $\epsilon$ computed over 1,000 test episodes for $\delta\in\{0.01,0.05,0.1\}$ across all instances (top), and for five randomly sampled instances with fixed $\delta=0.05$ (bottom). The results indicate substantial variability across instances. In some cases, $\epsilon$ is consistently positive, reflecting a clear predictive benefit; in others, it centers near zero or is negative, suggesting limited informativeness or even harmful effects on return prediction. Across instances, median values of $\epsilon$ remain close to zero for each evaluated $\delta$, exceeding zero only for $\delta=0.1$. These findings underscore the environment-dependent nature of $(\epsilon,\delta)$-informativeness.
@@ 7 CONCLUSION
We propose an informed asymmetric actor-critic method that generalizes the asymmetric approaches by allowing arbitrary privileged inputs in the critic without requiring full-state access, while preserving unbiased policy gradients. To guide the selection of such signals, we introduce two complementary informativeness criteria: a pre-training metric based on conditional dependence, and a post-hoc metric based on return prediction error. Our results show that privileged partial information can improve policy learning and value estimation, though its informativeness is task-dependent and not captured by return-based measures alone.
Future work may refine these criteria by exploring aspects such as signal complexity or direct policy impact, enabling better trade-offs between informativeness and model capacity for more robust and efficient learning under partial observability.
@@ Acknowledgments
Daniel Ebi gratefully acknowledges the financial support of the German Research Foundation (DFG) as part of the Research Training Group GRK 2153: Energy Status Data – Informatics Methods for its Collection, Analysis and Exploitation. Gaspard Lambrechts gratefully acknowledges the financial support of the Wallonia-Brussels Federation and the Fonds de la Recherche Scientifique (FNRS) for his FRIA grant. Additionally, this work was supported by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition.
@@ Appendix A AUXILIARY RESULTS
$Q^{\pi}(h,s,a)=R(s,a)+\gamma\,\\ E_{s^{\prime},o^{\prime}\mid s,a}\left[V^{\pi}(h^{\prime},s^{\prime})\right],$
with $s^{\prime}\sim T(s^{\prime}\mid s,a)$, $o^{\prime}\sim\widetilde{O}(o^{\prime}\mid s^{\prime})$, $i^{\prime}=s^{\prime}$, and $h^{\prime}$ denoting the updated history resulting from appending action $a$ and observation $o^{\prime}$ to $h$.
@@
$\nabla_{\theta}^{\text{AC}}J(\pi_{\theta})=\\ E\left[\sum_{t=0}^{\infty}\gamma^{t}\,Q^{\pi}(h_{t},s_{t},a_{t})\,\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid h_{t})\right].$
Following Lemma 4.1-4.3, this formulation recovers an alternative asymmetric policy gradient estimator that is equivalent to the standard policy gradient:

### v2

@@ 5 Experiments
We evaluate our informed asymmetric actor-critic framework along two complementary dimensions. First, we assess its learning performance on standard POMDP benchmarks against symmetric and asymmetric actor-critic baselines. Second, we validate the presented informativeness criteria in controlled informed POMDPs, where privileged signals can be systematically varied. We release our source code and experiments on GitHub [^1] https://github.com/EbiDa/informed-asymmetric-a2c to ensure reproducibility.
5.1 Benchmark Performance
Environments.
We use six benchmark navigation tasks from prior work [4] to evaluate the learning performance of our method: Heaven-Hell-3, Shopping-5, Car-Flag, Cleaner, Memory-Four-Rooms-7x7, and Memory-Four-Rooms-9x9. Each environment is formulated as a POMDP, and we define task-specific privileged signal accessible only to the critic. In Heaven-Hell-3 and Shopping-5, the privileged signal corresponds to the Manhattan distance to the target. In Car-Flag, the critic additionally observes the agent’s velocity. For the other tasks, the privileged input consists of an expanded spatial observation of the agent’s surroundings. See Appendix C.1 for details.
Baselines.
We compare our informed asymmetric actor-critic method, informed-asym-A2C, against three advantage actor-critic (A2C) variants: (1) A2C, a symmetric approach using a history-based critic $\hat{V}(h)$; (2) asym-A2C-s, an asymmetric variant with a state-based critic $\hat{V}(s)$; and (3) asym-A2C-hs, an asymmetric variant with a history-state critic $\hat{V}(h,s)$. All policy networks receive identical inputs and never access privileged information. Model architectures and hyperparameters follow Baisero and Amato [4] to ensure a fair comparison and are detailed in Appendix D.1.
@@ 5.1 Benchmark Performance
Results.
@@
Figure 1 reports the learning curves across all six benchmark tasks, showing episodic returns averaged over the last 100 episodes and aggregated across 20 independent runs. Across environments, informed-asym-A2C consistently improves sample efficiency and training stability relative to A2C. In Heaven-Hell-3, informed-asym-A2C exhibits strong performance relative to A2C and asym-A2C-s, though it is slightly outperformed by asym-A2C-hs. In Shopping-5, the informed asymmetric actor-critic converges faster than asym-A2C-hs and achieves comparable asymptotic return. In Car-Flag, it outperforms all baselines in both convergence speed and asymptotic return. In Cleaner, asym-A2C-hs achieves marginally higher returns, but informed-asym-A2C converges at a similar rate with greater stability than A2C, which suffers a performance drop after 2.5 million steps. In the Memory-Four-Rooms tasks, informed-asym-A2C substantially outperforms both asymmetric baselines, particularly asym-A2C-s.
Overall, informed-asym-A2C matches or surpasses the performance of asym-A2C-hs and/or achieves faster convergence in most tasks, demonstrating that appropriately structured privileged signals can match or outperform full-state informed critics while relying on strictly less additional information.
5.2 Validation of Informativeness Criteria
To empirically validate the proposed informativeness criteria in a controlled setting, we consider a distribution of synthetic informed POMDPs with finite state space ($|\mathcal{S}|=20$), discrete action space ($|\mathcal{A}|=4$), and continuous observation and privileged information spaces. Transition dynamics are generated following François-Lavet et al. [15] by sparsifying state-action transitions and normalizing to obtain valid probability distributions. Each state is associated with a latent Gaussian feature vector ($s_{t}\in\mathbb{R}^{5}$), and rewards are defined as linear functions of these features with weights $w_{r}$ sampled uniformly from [1,1][-1,1]. Privileged information is constructed by masking subsets of state features, while observations correspond to noisy subsets of this privileged signal. See Appendix C.2 for details on the environments.
Residual informativeness.
For the residual-based criterion, we estimate dependence between return-prediction residuals and privileged-signal residuals via the Hilbert-Schmidt Independence Criterion (HSIC) [19] with Gaussian RBF kernels, where bandwidths are chosen via the median heuristic and a Nyström approximation [28] with 512 landmarks is applied for efficiency. We collect 250 episodes of length 25 under a random policy and perform episode-level permutation testing with B=1, [^0] B=1{,}000 resamples. Conditional expectations are estimated via cross-fitting with a 100-tree random forest using 5-fold cross-validation. Observation-action histories are encoded with an RNN of hidden size 64, trained with temporal-difference learning on 100 training episodes.
Post-hoc informativeness.
For the post-hoc criterion, we train symmetric and informed critics with temporal-difference learning using 5-fold cross-validation on the same environments for 2,500 episodes of length 25. Histories are encoded with an RNN of hidden size 64, and the resulting fixed-length representation is concatenated with the privileged signal and fed into a linear output layer for value prediction. We evaluate episode-level gains in value prediction accuracy across held-out folds, with $\epsilon=0$.
Results.
Table 1 reports results for multiple privileged signal generators under both criteria across ten independent runs on a randomly sampled environment. In this setting, the agent observes noiseless $o_{t}=(s_{t}^{1},s_{t}^{2})$, while only a subset of state components ($s_{t}^{4},s_{t}^{5}$) are relevant for the rewards, i.e., $w_{r}~=~[0.0001,0.0001,-0.0001,-1.0,1.0]$. Privileged signals that include these return-relevant components consistently exhibit stronger residual dependence and positive prediction gains. In contrast, signals containing only redundant or irrelevant ($s_{t}^{3}$) components yield weaker statistical evidence under both criteria. Interestingly, full-state access does not always achieve the best scores, highlighting that additional features can be uninformative when misaligned with return prediction. These results demonstrate that the two criteria provide complementary insights into signal informativeness, and that carefully selecting signals that capture return-relevant state components is most effective. Ablations and further results are provided in Appendix E.
Table 1: Comparison of privileged signal generators using residual-based and post-hoc prediction informativeness criteria across ten independent runs on a randomly sampled informed POMDP environment with reward weights $w_{r}~=~[0.0001,0.0001,-0.0001,-1.0,1.0]$. All values are reported as mean $\pm$ standard deviation.
Privileged signal $i_{t}$ Residual Informativeness Prediction Informativeness
$\rho_{\text{obs}}$ ($\bar{x}\pm\hat{\sigma}$) $p$-value ($\bar{x}\pm\hat{\sigma}$) $L_{\tau}$ ($\bar{x}\pm\hat{\sigma}$) $p$-value ($\bar{x}\pm\hat{\sigma}$)
$i_{t}=[s_{t}^{1},s_{t}^{2}]$ 3.0e-05 $\pm$ 8.1e-06 0.438 $\pm$ 0.290 -1.4e-02 $\pm$ 0.014 0.801 $\pm$ 0.330
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{3}]$ 5.5e-05 $\pm$ 1.9e-05 0.170 $\pm$ 0.169 0.007 $\pm$ 0.016 0.397 $\pm$ 0.338
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{4}]$ 7.6e-05 $\pm$ 3.2e-05 0.119 $\pm$ 0.262 0.035 $\pm$ 0.018 0.043 $\pm$ 0.077
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{5}]$ 5.0e-05 $\pm$ 1.5e-05 0.191 $\pm$ 0.228 0.018 $\pm$ 0.019 0.232 $\pm$ 0.248
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{3},s_{t}^{4}]$ 7.4e-05 $\pm$ 2.4e-05 0.068 $\pm$ 0.116 0.046 $\pm$ 0.018 0.009 $\pm$ 0.016
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{3},s_{t}^{5}]$ 5.8e-05 $\pm$ 1.5e-05 0.106 $\pm$ 0.141 0.026 $\pm$ 0.019 0.120 $\pm$ 0.152
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{4},s_{t}^{5}]$ 7.6e-05 $\pm$ 2.1e-05 0.035 $\pm$ 0.051 0.064 $\pm$ 0.020 0.003 $\pm$ 0.006
$i_{t}=[s_{t}^{1},s_{t}^{2},s_{t}^{3},s_{t}^{4},s_{t}^{5}]$ 7.0e-05 $\pm$ 1.8e-05 0.038 $\pm$ 0.051 0.056 $\pm$ 0.032 0.072 $\pm$ 0.131
@@ 6 Conclusion
We study asymmetric actor-critic learning under training-execution asymmetry and introduce the informed asymmetric actor-critic framework, which allows the critic to condition on state-dependent privileged signals unavailable at deployment while preserving unbiased policy gradients. This generalization moves asymmetric learning beyond full-state critics and shifts the focus from state access to identifying informative training-time signals. We propose two complementary criteria to assess informativeness before training and post hoc, and empirically demonstrate that appropriately chosen privileged signals can substantially improve learning and value estimation in POMDPs.
Future work may extend these criteria to account for signal complexity or direct policy effects, enabling better trade-offs between informativeness and model capacity for more robust and efficient learning under partial observability.
@@ Acknowledgments
Daniel Ebi gratefully acknowledges the financial support of the German Research Foundation (DFG) as part of the Research Training Group GRK 2153: Energy Status Data – Informatics Methods for its Collection, Analysis and Exploitation. This work was carried out while Gaspard Lambrechts was a postdoctoral researcher at the University of Liège, supported by the Fund for Scientific Research (FNRS) from the Wallonia-Brussels Federation in Belgium. Additionally, this work was supported by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Auxiliary Results
$Q^{\pi}(h,s,a)=R(s,a)+\gamma\,\\ E_{s^{\prime},o^{\prime}}\left[V^{\pi}(h^{\prime},s^{\prime})|s,a\right],$
with $s^{\prime}\sim T(s^{\prime}|s,a)$, $o^{\prime}\sim\widetilde{O}(o^{\prime}|s^{\prime})$, $i^{\prime}=s^{\prime}$, and $h^{\prime}$ denoting the updated history resulting from appending action $a$ and observation $o^{\prime}$ to $h$.
@@
$\nabla_{\theta}^{\text{AC}}J(\pi_{\theta})=\\ E\left[\sum_{t=0}^{\infty}\gamma^{t}\,Q^{\pi}(h_{t},s_{t},a_{t})\,\nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\right].$
Following Lemma 3.1 and 3.3, this formulation recovers an alternative asymmetric policy gradient estimator that is equivalent to the standard policy gradient: